# Volumetric Functional Maps

### v2

@@ Abstract.
The computation of volumetric correspondences between 3D shapes is a prominent tool for medical and industrial applications. In this work, we pave the way for spectral volume mapping, extending for the first time the functional maps framework from the surface to the volumetric setting. We show that the eigenfunctions of the volumetric Laplace operator define a functional space that is suitable for high-quality signal transfer. We also experiment with various techniques that edit this functional space, porting them to volume domains. We validate our method on novel volumetric datasets and on tetrahedralizations of well established surface datasets, also showcasing practical applications involving both discrete and continuous signal mapping, for segmentation transfer, mesh connectivity transfer and solid texturing. Last but not least, we show that considering the volumetric spectrum greatly improves the accuracy for classical shape matching tasks among surfaces, consistently outperforming existing surface-only spectral methods.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: TOGjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
@@
Figure 1. A visual representation of our pipeline. The eigenfunctions of the Laplace-Beltrami operator for tetrahedral meshes (left) are used to compute a functional map among volumes (middle). The basis alignment can be applied for several tasks, such as computation of volumetric correspondences, generation of piece-wise linear simplicial maps, and volumetric segmentation transfer (right).
1. Introduction
@@
The computation of correspondences between volumetric objects plays a central role for many critical applications. Volume mappings enable non-invasive medical diagnosis, by warping the digital copy of a human organ into its canonical position without extracting it from the patient’s body (Abulnaga et al., 2021). They are also important for hexmeshing (Pietroni et al., 2022), statistical shape analysis (Huang et al., 2019) and for Industry 4.0, where physical simulations computed on a predictive digital twin must be aligned with the other digital views of the same real environment for optimized decision making and failure avoidance (Rasheed et al., 2020).
@@
Maps between 3D objects are often represented in a computer either by exploiting one-to-one per vertex correspondences between two simplicial meshes with same connectivity, defining a so called piece-wise map (Section 2.1), or putting into correspondence two sets of basis functions that allow to express the same signal in two alternative domains, as done by spectral approaches for surface meshes (Ovsjanikov et al., 2012). Both piece-wise and spectral approaches have been extensively studied in the surface setting and can nowadays be considered rather mature, as proved by the numerous academic and commercial tools that incorporate these technologies, such as Blender, Maya and 3DS Max. Conversely, the equivalent task of computing volumetric correspondences between solid objects largely remains a scientific open problem (Pietroni et al., 2022; Nigolian et al., 2023, §8.2).
@@
In recent years various attempts have been made to extend piece-wise linear approaches to the volume setting. This has proved to be an extremely challenging task, for which no fully robust and efficient methods have been devised yet (Section 2.1). Conversely, to our knowledge no attempt has been made so far to extend functional mapping from the surface to the volume setting (Section 2.2).
In this article, for the first time, we propose this extension and show that, conversely from the piece-wise linear setting, functional mapping naturally scales to volumes with minimal effort, providing a theoretically sound and practically useful platform for the computation of volumetric correspondences between solid shapes.
@@
Using the spectrum of the volumetric Laplacian, we define a functional space that is completely agnostic with respect to the underlying discrete mesh, and we use it to create correspondences between solid meshes with different density and connectivity. Interestingly, in the paper we show that many recent developments used to improve efficiency by reducing the size of the functional space without sacrificing accuracy (Melzi et al., 2019b; Maggioli et al., 2021) can also be exploited in the volume setting, permitting the finding of an optimal balance between map quality and performance.
@@
We validate our volumetric pipeline (showcased in Figure 1) on a variety of datasets, including tetrahedralizations of prominent reference benchmarks widely used in the functional mapping literature. Besides measuring the accuracy and performances of our tool on ground truth data (Section 5), we showcase a variety of practical applications, including connectivity transfer of a given tetrahedral mesh to a target domain to generate a piece-wise linear map (Figure 7), segmentation transfer from a statistical medical shape to acquired organs (Figure 10) and solid texturing (Figure 6).
@@
2. Related Works
Here, we discuss the main approaches for the computation of volume maps with the piece-wise linear paradigm and surface correspondence with the functional maps framework. In doing this, we highlight the existing attempts to scale from the surface to the volume setting (or lack thereof) for each class of algorithms.
2.1. Piece-wise Maps
@@
Piece-wise maps define correspondences between two simplicial meshes sharing the same amount of vertices and same connectivity. Starting from explicit one-to-one per vertex correspondences the map is implicitly extended inside mesh elements by means of linear barycentric interpolation. This operation is uniquely defined for simplices in any dimension (Hormann, 2014) and can therefore be used for both surface (triangle) and volume (tetrahedral) meshes.
A variety of algorithms ensure the ability to generate piece-wise linear surface maps to elementary domains such as triangles (Finnendahl et al., 2023), convex polygons (Tutte, 1963; Floater, 1997; Livesu, 2024b; Shen et al., 2019), star-shaped polygons (Livesu, 2024a) and spheres (Gotsman et al., 2003; Praun and Hoppe, 2003). Mappings to these basic shapes (or to assemblies of them) are the main ingredient to generate higher level maps between general shapes, which are typically obtained via composition, overlaying two surface maps one onto the other to define inter-surface correspondences (Schmidt et al., 2019; Schreiner et al., 2004; Weber and Zorin, 2014; Lee et al., 1998; Khodakovsky et al., 2003; Praun et al., 2001; Schmidt et al., 2020; Livesu, 2021; Garner et al., 2005; Shi et al., 2016; Aigerman and Lipman, 2016; Li et al., 2008; Schmidt et al., 2023; Kraevoy and Sheffer, 2004).
@@
Lifting the same idea to the volume setting is possible in principle but extremely challenging in practice (Livesu, 2020a; Cherchi and Livesu, 2023; Nigolian et al., 2024). Even constructing an injective map to a basic convex polyhedron has proved to be a surprisingly difficult task. Surface algorithms such as the Tutte embedding are known to fail on general volume meshes (Alexa, 2023; Chilakamarri et al., 1995). More recent methods, such as (Livesu, 2024b, a; Finnendahl et al., 2023) may potentially scale to volumes but concrete extensions do not exist yet. To date, the only existing algorithms for robust volume embedding are (Nigolian et al., 2023, 2024; Hinderink and Campen, 2023; Campen et al., 2016). All of them are extremely inefficient both in terms of running times and memory consumption, mainly due to the need of using exact numerical constructions and exponential mesh refinement. They are therefore unsuitable for practical applications.
@@
A less robust but more practical alternative to the map composition pipeline consists in initializing a non injective volume map directly, with Tutte 3D or some other non robust method, and then restore the correct orientation for elements that flip or vanish through the map via vertex relocation. In the volume setting, orientation (injectivity) constraints translate to sign constraints on the determinant of the $3\times 3$ Jacobian matrix associated to each per element transformation, leading to an optimization problem subject to cubic constraints which is numerically intractable. Most of the research in the field focuses on how to formulate (and relax) the orientation constraints to make the problem tractable. Removing inverted elements from a given mesh is also a well known problem in computer simulation, where it is often referred to as mesh untangling (Knupp, 2001; Pietroni et al., 2022). Note that in the general case mesh untangling is an ill-posed problem, because it may not admit a valid solution for a fixed mesh connectivity. Additional degrees of freedom must be inserted via mesh refinement to open the feasible region (see (Livesu, 2020b, Fig.1) for a 2D example). Existing volume approaches employ minimization of flip-preventing energies (Du et al., 2020; Garanzha et al., 2021; Abulnaga et al., 2023), operate on convex subsets of the feasible region (Kovalsky et al., 2014) or project onto the space of bounded distortion mappings (Aigerman and Lipman, 2013). Due to the non linearity of the problem and the non existence of a feasible starting solution, none of these methods guarantees the correctness of the result. Multiple failure cases have been reported by testing the most prominent existing algorithms on available datasets (Nigolian et al., 2023, §6.3). In Section 5.2 we show that bootstrapping this pipeline with our functional mapping dramatically increases the success rate of state of the art volume untangling algorithms.
2.2. Functional Maps
@@
One of the most fundamental topics in shape analysis is the shape matching task. Given a pair of discretized surfaces of 3D objects, shape matching estimates a meaningful pointwise correspondence between them. If the shapes are rigid, we refer to this task as rigid shape matching. Otherwise, we refer to it as deformable shape matching or, for brevity, shape matching, which is more general and usually more challenging to solve. Many applications are related to this task, from statistical shape analysis (Bogo et al., 2014) and medical imaging (Magnet et al., 2023) to deformation transfer (Sumner and Popović, 2004) among many others. For this reason, many researchers target this task (Sahillioğlu, 2020; Deng et al., 2022). Among the different approaches, the functional maps framework (Ovsjanikov et al., 2012, 2016) has attracted significant attention. This framework proposes to solve the shape matching problem by estimating a correspondence between functions defined on the surfaces instead of focusing on correspondence among points which are hard to optimize. Furthermore, exploiting a basis for the space of functions defined on the surfaces, for which a standard choice arises from the extension of Fourier theory to non-Euclidean domains (Taubin, 1995; Lévy and Zhang, 2010), the correspondence among functions can be compactly encoded in a matrix with dimensions equal to the size of the selected basis. Following this direction, a variety of works tried to improve the framework by defining additional functional constraints (Nogneng and Ovsjanikov, 2017; Ren et al., 2018; Donati et al., 2022), proposing alternative bases (Melzi et al., 2018; Nogneng et al., 2018; Maggioli et al., 2021; Marin et al., 2019, 2020), or defining an alternative procedure to extract the correspondence from a given functional map (Rodola et al., 2017; Melzi et al., 2019b; Ren et al., 2020, 2021; Ezuz and Ben-Chen, 2017; Pai et al., 2021). Recently, by proposing two alternatives for scaling the functional maps approach to high-resolution meshes (Magnet and Ovsjanikov, 2023; Maggioli et al., 2025) addressed the issue of scalability of the functional maps framework. The functional maps framework additionally gives rise to a family of data-driven approaches that exploit machine learning techniques to solve the shape matching problem exploiting the functional representation as a prior (Litany et al., 2017; Donati et al., 2020; Sharp et al., 2022; Cao et al., 2024, 2023; Attaiki and Ovsjanikov, 2024). The ones listed above are just a subset of the literature related to the functional maps framework. Its impact in computer graphics and related fields has been impressive, and many applications have been addressed exploring this technique, from statistical shape analysis in medicine (Maccarone et al., 2024) to shape deformation (Sundararaman et al., 2024). Despite its success, to the best of our knowledge, this framework has never been applied to volumetric data, leaving its potential impact entirely unexplored in volumetric data processing and related tasks.
3. Background and Notation
@@
Scalar functions $f:V_{\mathcal{M}}\to\mathbb{R}$ over a volume mesh $\mathcal{M}$ are discretized as vectors $\bm{f}\in\mathbb{R}^{|V_{\mathcal{M}}|}$. Given a scalar function $f:V_{\mathcal{M}}\to\mathbb{R}$ over a volume mesh $\mathcal{M}$, its restriction $f_{\partial}:V_{\partial\mathcal{M}}\to\mathbb{R}$ to the external surface $\partial\mathcal{M}$ is discretized as a vector $\bm{f}_{\partial}\in\mathbb{R}^{|V_{\partial\mathcal{M}}|}$.
@@
Given a tetrahedral (or triangular) mesh $\mathcal{M}$, and given two vertices $v_{i},v_{j}\in\mathcal{M}$, we refer to the geodesic distance between $v_{i}$ and $v_{j}$ over $\mathcal{M}$ as $d_{\mathcal{M}}(v_{i},v_{j})$. There exist several algorithms to estimate such a distance on discrete meshes (Crane et al., 2020). For the sake of simplicity, in the remainder of the paper we will always measure $d_{\mathcal{M}}(v_{i},v_{j})$ using Dijkstra’s algoritm (1959) on the edge connectivity of $\mathcal{M}$.
Shape Matching and Correspondences
@@
Given two surface meshes $\mathcal{M}=(V_{\mathcal{M}},F_{\mathcal{M}})$ and $\mathcal{N}=(V_{\mathcal{N}},F_{\mathcal{N}})$, the problem of shape matching is the problem of finding a correspondence $\pi:V_{\mathcal{M}}\to V_{\mathcal{N}}$ that maps vertices of $\mathcal{M}$ into vertices of $\mathcal{N}$ (Loncaric, 1998). In most cases, the correspondence $\pi$ has to satisfy semantic constraints or has to be an isometry (Sahillioğlu, 2020; Deng et al., 2022). An isometry is a correspondence $\pi$ that preserves the geodesic distances between pairs of vertices, namely
(1) $\forall v_{i},v_{j}\in V_{\mathcal{M}},\quad d_{\mathcal{M}}(v_{i},v_{j})=d_{\mathcal{N}}(\pi(v_{i}),\pi(v_{j}))\,.$
Functional Maps
@@
The problem of finding an isometric correspondence, in its general setting, is NP-hard (Benkner et al., 2021), and thus many solutions have been developed for computing an approximation. Among these, the functional maps approach has grown popular due to its high efficiency and the smoothness of the resulting correspondences (Ovsjanikov et al., 2012).
@@
Instead of searching directly for a correspondence $\pi$ between two surfaces $\mathcal{M}$ and $\mathcal{N}$, the functional maps approach aims at finding a linear map $T_{\pi}:\mathcal{F}(\mathcal{N})\to\mathcal{F}(\mathcal{M})$, induced by $\pi$, from the space $\mathcal{F}(\mathcal{N})$ of real-valued functions over $\mathcal{N}$ to the space $\mathcal{F}(\mathcal{M})$ of real-valued functions over $\mathcal{M}$. Given any function $f:V_{\mathcal{N}}\to\mathbb{R}$, the map $T_{\pi}$ gives the corresponding function $T_{\pi}(f):V_{\mathcal{M}}\to\mathbb{R}$ that maps $v_{i}\mapsto f(\pi(x))$, for each $v_{i}\in V_{\mathcal{M}}$.
@@
Given two sets of (orthogonal) basis functions $\Phi_{\mathcal{M}},\Phi_{\mathcal{N}}$, respectively for the functional space $\mathcal{F}(\mathcal{M}),\mathcal{F}(\mathcal{N})$, and a function $f:V_{\mathcal{N}}\to\mathbb{R}$, the function $f$ can be represented as a linear combination of the basis functions $\Phi_{\mathcal{N}}$ as
(2) $f=\sum_{i=0}^{\infty}\alpha_{i}\phi_{\mathcal{N}}^{(i)}=\sum_{i=0}^{\infty}\langle f,\phi_{\mathcal{N}}^{(i)}\rangle_{\mathcal{N}}\phi_{\mathcal{N}}^{(i)}\,,$
@@
where $\phi_{\mathcal{N}}^{(i)}$ is the $i$-th basis function in $\Phi_{\mathcal{N}}$ and $\langle\cdot,\cdot\rangle_{\mathcal{N}}$ denotes the inner product between functions over $\mathcal{N}$. The coefficients $\alpha_{i}$ thus denotes the projection of $f$ onto the basis functions $\phi_{\mathcal{N}}^{(i)}$. A usual choice for the bases $\Phi_{\mathcal{M}},\Phi_{\mathcal{N}}$ are the eigenfunctions of the Laplace-Beltrami operator on $\mathcal{M}$ and $\mathcal{N}$, respectively.
Since $T_{\pi}$ is linear, the mapping of $f$ onto $\mathcal{M}$ can then be computed as
(3) $T_{\pi}(f)=\sum_{i=0}^{\infty}\alpha_{i}T_{\pi}(\phi_{\mathcal{N}}^{(i)})=\sum_{i,j=0}^{\infty}\alpha_{i}c_{i,j}\phi_{\mathcal{M}}^{(j)}\,,$
@@
(4) $T_{\pi}(f)\approx\Phi_{\mathcal{M}}\mathbf{C}\Phi_{\mathcal{N}}^{\dagger}\bm{f}\,,$
where denotes the Moore-Penrose pseudo-inverse.
@@
Notably, $T_{\pi}(\Phi_{\mathcal{N}})\approx\Phi_{\mathcal{M}}\mathbf{C}$, and thus by knowing the matrix $\mathbf{C}$ it is possible to extract the unknown correspondence $\pi$ via a nearest-neighbor search $\mathrm{NN}(\Phi_{\mathcal{N}},\Phi_{\mathcal{M}}\mathbf{C})$.
4. Method

### v3

@@ Abstract
Computing volumetric correspondences between 3D shapes is a prominent tool for medical and industrial applications. In this work, we pave the way for spectral volume mapping, extending for the first time the surface-based functional maps framework. We show that the eigenfunctions of the volumetric Laplace operator define a functional space that is suitable for high-quality signal transfer. We also experiment with various techniques that edit this functional space, porting them to volume domains. We validate our method on novel volumetric datasets and on tetrahedralizations of well established surface datasets, also showcasing practical applications involving both discrete and continuous signal mapping, for segmentation transfer, mesh connectivity transfer and solid texturing. Finally, we show that the volumetric spectrum greatly improves the accuracy for classical shape matching tasks among surfaces, consistently outperforming surface-only spectral methods.
@@
Figure 1: Visual representation of our pipeline. The eigenfunctions of the LBO for volume meshes (left) are used to compute a volumetric functional map (middle). Basis alignment is exploited for several tasks: volumetric correspondences, piece-wise linear maps, and volumetric segmentation transfer (right).
1 Introduction
@@
The computation of volumetric correspondences plays a central role for many applications. Volume mappings enable non-invasive medical diagnosis, by warping the digital copy of a human organ into its canonical position without extracting it from the patient’s body [2]. They are also important for meshing [91], statistical shape analysis [44] and Industry 4.0, where physical simulations computed on predictive digital twins must be aligned with other digital views for optimized decision making and failure avoidance [96].
@@
Maps between 3D objects are often represented either by exploiting one-to-one per vertex correspondences between two simplicial meshes with same connectivity, defining a so called piece-wise map (Sec. 2.1), or putting into correspondence two sets of basis functions to express the same signal in two alternative domains, as done by spectral approaches for surface meshes [87]. Both piece-wise and spectral approaches have been extensively studied in the surface setting and can nowadays be considered rather mature, as proved by numerous academic and commercial tools that incorporate them. Conversely, the equivalent task of computing volumetric correspondences largely remains a scientific open problem [91, §8.2]. It is worth noticing that the LDDMM framework has been successfully applied to match dense volumetric data (e.g., MRI/CRT scans), although diffeomorphic maps are limited to domains with grid-like structures and hardly generalize to more flexible representations like tetrahedral meshes[16, 112, 92, 17].
@@
In recent years many attempts have been made to extend piece-wise linear approaches to volumes. This task has proved to be extremely challenging and no fully robust and efficient methods have been devised yet (Sec. 2.1). To our knowledge no attempt has been made to extend functional mapping to volumes yet (Sec. 2.2). We propose it for the first time and show that, conversely from the piece-wise linear setting, functional mapping naturally scales to volumes, providing a theoretically sound and practically useful platform for the computation of voulmetric correspondences.
@@
Using the spectrum of the volumetric Laplacian, we define a functional space that is completely agnostic of the underlying discrete mesh, and we use it to create correspondences between solid meshes with different density and connectivity. We also show that many recent developments used to improve the functional maps framework [79, 70] can be exploited on volumes too, allowing to find an optimal balance between map quality and performance.
@@
We validate our volumetric pipeline (Fig. 1) on a variety of datasets, including tetrahedralizations of prominent reference benchmarks widely used for shape correspondence tasks. Besides assessing the accuracy and performances of our tool on the ground truth (Sec. 5), we showcase various practical applications, including connectivity transfer of a given volume mesh to a target domain (Fig. 7), segmentation transfer from a statistical medical shape to acquired organs (Fig. 8) and solid texturing (Fig. 5).
@@
2 Related Works
We list the main approaches to compute piece-wise maps and functional maps highlighting, for both classes of methods, existing attempts (or lack thereof) to scale to volumes.
2.1 Piece-wise Maps
@@
Piece-wise maps define correspondences between two meshes by starting from explicit per vertex correspondences and extending the map inside mesh elements by means of barycentric interpolation. This operation is uniquely defined for simplices in any dimension [42], therefore it works for both surface (triangle) and volume (tetrahedral) meshes.
A variety of algorithms can robustly generate piece-wise surface maps to elementary domains such as triangles [33], convex polygons [118, 35, 65, 109], star-shaped polygons [64] and spheres [38, 94]. Maps between general shapes are then obtained via composition, using these elementary shapes as intermediate domains [104, 107, 120, 52, 46, 95, 105, 63, 37, 110, 6, 56, 106, 49].
@@
Lifting this idea to volumes is possible in principle but extremely challenging in practice. Even constructing an injective map to a convex polyhedron has proved to be surprisingly difficult [61, 18, 84]. Surface methods such as the Tutte embedding are known to fail on general volume meshes [8, 19]. To date, the only robust volume embedding algorithms are [83, 84, 41, 13], but all of them are unsuitable for real applications, due the massive use of exact numerical constructions and exponential mesh refinement [75].
@@
A less robust but more practical alternative consists in initializing a non injective map directly with a fast approximation algorithm, and then relocate the vertices to fix flipped or vanishing elements. This problem, often referred to as mesh untangling [47, 91], is ill-posed in general, as valid solutions may not exist for a fixed mesh connectivity [62]. Existing volume approaches operate on convex subsets of the feasible region [48], minimize flip-preventing energies [28, 36, 1], or project onto the space of bounded distortion mappings [5]. Due to the problem’s non linearity and the absence of a feasible starting solution, none of these methods guarantees the correctness of the result. Multiple failure cases have been reported by testing the most prominent existing algorithms on available datasets [83, §6.3]. In Sec. 5.2 we show that bootstrapping this pipeline with our functional mapping dramatically increases the success rate of state of the art volume untangling algorithms.
2.2 Functional Maps
@@
Given two discrete surfaces, shape matching aims to estimate a semantic point-wise correspondence between them. The more general (and challenging) scenario occurs when the two shapes are related by a non-rigid deformation. Many applications are related to this task, from statistical shape analysis [11] and medical imaging [71] to deformation transfer [114, 115], among many others [103, 22]. Among different approaches, functional mapping [87, 88] has attracted significant attention. This framework aims to solve the shape matching problem by estimating a correspondence between functions defined on the surfaces, rather than point-wise correspondences, which are hard to optimize for. Exploiting a basis for the space of functions, for which a standard choice arises from the extension of Fourier theory to non-Euclidean domains [116, 54], the correspondence among shapes can be compactly encoded in a matrix with dimensions equal to the basis size. Following this direction, a variety of works try to improve the framework by defining additional functional constraints [86, 100, 26], proposing alternative bases [80, 85, 70, 73, 74], or defining alternative procedures to extract point correspondences from a given functional map [102, 79, 98, 99, 32, 89]. Recently, two alternatives for scaling the functional maps approach to high-resolution meshes [72, 69] addressed the scalability issues of functional mapping. The functional maps framework additionally gives rise to a family of data-driven approaches that exploit machine learning techniques to solve the shape matching problem exploiting the functional representation as a prior [58, 27, 108, 14, 15, 9]. Despite its success, to the best of our knowledge, this framework has never been applied to volumes, leaving its potential impact fully unexplored in volume data processing.
3 Background and Notation
@@
A scalar function $f:V_{\mathcal{M}}\to\mathbb{R}$ over a volume mesh $\mathcal{M}$ is discretized as a vector $\bm{f}\in\mathbb{R}^{|V_{\mathcal{M}}|}$, and its restriction $f_{\partial}:V_{\partial\mathcal{M}}\to\mathbb{R}$ to the external surface $\partial\mathcal{M}$ is discretized as a vector $\bm{f}_{\partial}\in\mathbb{R}^{|V_{\partial\mathcal{M}}|}$. Given a tetrahedral (or triangular) mesh $\mathcal{M}$, and given two vertices $v_{i},v_{j}\in\mathcal{M}$, we refer to the geodesic distance between $v_{i}$ and $v_{j}$ over $\mathcal{M}$ as $d_{\mathcal{M}}(v_{i},v_{j})$. There exist several algorithms to estimate such a distance on discrete meshes [20]. For the sake of simplicity, in the remainder of the paper we will always approximate $d_{\mathcal{M}}(v_{i},v_{j})$ using Dijkstra’s algoritm [25] on the edge connectivity of $\mathcal{M}$.
Shape Matching and Correspondences.
@@
Given two surface meshes $\mathcal{M}=(V_{\mathcal{M}},F_{\mathcal{M}})$ and $\mathcal{N}=(V_{\mathcal{N}},F_{\mathcal{N}})$, the problem of shape matching is the problem of finding a correspondence $\pi:V_{\mathcal{M}}\to V_{\mathcal{N}}$ that maps vertices of $\mathcal{M}$ into vertices of $\mathcal{N}$ [66]. In most cases, the correspondence $\pi$ has to satisfy semantic constraints or has to be an isometry [103, 22], that is a correspondence $\pi$ preserving the geodesic distances between pairs of vertices. Namely
$\forall v_{i},v_{j}\in V_{\mathcal{M}},\quad d_{\mathcal{M}}(v_{i},v_{j})=d_{\mathcal{N}}(\pi(v_{i}),\pi(v_{j}))\,.$ (1)
Functional Maps.
@@
The problem of finding an isometric correspondence, in its general setting, is NP-hard [10]. Many solutions have been developed for computing an approximation. Among these, the functional maps approach has grown popular due to its high efficiency and the smoothness of the resulting correspondences [87].
@@
Instead of searching directly for a correspondence $\pi$ between two surfaces $\mathcal{M}$ and $\mathcal{N}$, the functional maps approach aims at finding a linear map $T_{\pi}:\mathcal{F}(\mathcal{N})\to\mathcal{F}(\mathcal{M})$, induced by $\pi$, from the space $\mathcal{F}(\mathcal{N})$ of real-valued functions over $\mathcal{N}$ to the space $\mathcal{F}(\mathcal{M})$ of real-valued functions over $\mathcal{M}$. Given any function $f:V_{\mathcal{N}}\to\mathbb{R}$, the map $T_{\pi}$ gives the corresponding function $T_{\pi}(f):V_{\mathcal{M}}\to\mathbb{R}$ that maps $v_{i}\mapsto f(\pi(v_{i}))$, for each $v_{i}\in V_{\mathcal{M}}$.
@@
Given two sets of (orthogonal) basis functions $\Phi_{\mathcal{M}},\Phi_{\mathcal{N}}$ for the spaces $\mathcal{F}(\mathcal{M}),\mathcal{F}(\mathcal{N})$, respectively, a function $f:V_{\mathcal{N}}\to\mathbb{R}$ can be expressed as a linear combination of $\Phi_{\mathcal{N}}$
$f=\sum_{i=0}^{\infty}\alpha_{i}\phi_{\mathcal{N}}^{(i)}=\sum_{i=0}^{\infty}\langle f,\phi_{\mathcal{N}}^{(i)}\rangle_{\mathcal{N}}\phi_{\mathcal{N}}^{(i)}\,$, (2)
@@
where $\phi_{\mathcal{N}}^{(i)}$ is the $i$-th basis function in $\Phi_{\mathcal{N}}$ and $\langle\cdot,\cdot\rangle_{\mathcal{N}}$ denotes the inner product between functions over $\mathcal{N}$. The coefficients $\alpha_{i}$ denotes the projection of $f$ onto the basis functions $\phi_{\mathcal{N}}^{(i)}$. Usually, $\Phi_{\mathcal{M}},\Phi_{\mathcal{N}}$ are chosen to be the eigenfunctions of the Laplace-Beltrami operator on $\mathcal{M}$ and $\mathcal{N}$. Since $T_{\pi}$ is linear, the mapping of $f$ onto $\mathcal{M}$ is given by
$T_{\pi}(f)=\sum_{i=0}^{\infty}\alpha_{i}T_{\pi}(\phi_{\mathcal{N}}^{(i)})=\sum_{i,j=0}^{\infty}\alpha_{i}c_{i,j}\phi_{\mathcal{M}}^{(j)}\,$, (3)
@@
$T_{\pi}(f)\approx\Phi_{\mathcal{M}}\mathbf{C}\Phi_{\mathcal{N}}^{\dagger}\bm{f}\,$, (4)
@@
where denotes the Moore-Penrose pseudo-inverse. Notably, $T_{\pi}(\Phi_{\mathcal{N}})\approx\Phi_{\mathcal{M}}\mathbf{C}$, thus by knowing the matrix $\mathbf{C}$ it is possible to extract the unknown correspondence $\pi$ via a nearest-neighbor search, $\mathrm{NN}(\Phi_{\mathcal{N}},\Phi_{\mathcal{M}}\mathbf{C})$.
4 Method